# Prompts Are Programs Too! Understanding How Developers Build Software Containing Prompts

### v1

@@ 5. Prompt Programming Process
We found that the prompt programming process exhibited significant differences across each of the software development activities (see Figure 2). We identified six types of activities in prompt programming: requirements, design, implementation, debugging, data curation, and evaluation. We first present these results (see Section 5.1) and conclude with its implications (see Section 5.2).
5.1. Results
5.1.1. Requirements
@@
Requirements underpin the construction of prompt programs. Similar to prior work (Furmakiewicz et al., 2024; Parnin et al., 2023), participants noted the existence of functional and non-functional requirements, such as usability (P13, P15), reliability (P7, P19), latency (P4, P11, P15, P19), safety (P7, P10, P13, P19), privacy (P4, P13, P14), availability (P4), and cost (P5, P8, P10, P11, P14, P16, P19, P20). We observed that companies found safety and privacy especially important, as noted in the literature (Hassan et al., 2024; Patel et al., 2024).
Assumptions of the requirements must be explicitly stated
@@
At times, the FM did not perform the correct action not because of an FM error, but due to an under-specified or conflicting requirement in the prompt (P3, P4, P5, P6, P8, P12, P19). Thus, prompt programming required explicitly stating assumptions: “When you’re interacting with a person, you make all these assumptions, but everybody can read between the lines… The more you design prompts, the more you realize that it’s not the same as communicating with another person. The model needs more specific instructions” (P3). However, predicting which assumptions to elaborate on was challenging since “it’s hard to tell when the model could go wrong” (P3): “I think observation is like 80% [of the requirements]. I knew what I wanted 20% of the time” (P8). As a result, “capturing all the rules…is very difficult in the start, and just generalizing over all the use cases there can be very difficult” (P19).
Requirements can evolve as the capabilities of the FM are discovered
@@
Similar to building ML-enabled systems (Wan et al., 2019), requirements of the prompt could change as the capability of the FM on the task was discovered. Changed requirements included making new trade-offs between non-functional requirements, such as safety versus quality (Furmakiewicz et al., 2024) (P7): “Our requirements did change as to how the product was perceived [by users]… So, striving for that balance between the response quality and the latency induced quite a few iterations just to see how much room do we have to play with” (P15). Additionally, participants noted defining new features based on observing how the FM performed on the task (P2, P4, P13, P18, P20): “I do add in requirements as I iterate. ‘Oh, I didn’t know that you would be able to do this. Let’s see how much I could push this functionality”’ (P2).
5.1.2. Design
Complex or lengthier prompts can be decomposed into into multiple components.
Prompt programs can be composed and decomposed.
@@ 5.1.2. Design
Participants decomposed prompt programs into smaller components based on the complexity of the prompt program. For a single prompt, participants described decomposing the prompt into sub-components, such as into subsections (P1, P2, P5, P6, P7, P8, P13, P20). One participant noted “it is important to structure the prompt in a way that makes sense to you. Otherwise, you’re not able to maintain it going forward” (P13).
@@
Prompts can be composed or “chained” together for greater capabilities (Liu et al., 2023; Wu et al., 2022a, b; Arawjo et al., 2023). Participants chained multiple prompts that depended on each other (see Table 1) for complex tasks: “We know what the complex task looks like, and then our job is to break that down…Evaluating each step independently of the others can help for complex tasks” (P9). Each prompt had its own responsibility: “if you are creating a robot, the motion of the hand in any scenario could be a responsibility to one agent. The movement of the legs could be given as a responsibility to another agent… Sometimes, they have to communicate with each other to perform a synchronous action” (P19). Similar to pure code systems, prompts acted as a single module with its own concern (Dijkstra, 1982) which could be composed into a more complex system. Yet, composing multiple prompt programs posed challenges (see Section 4.1.3), causing coupling as the prompts were modified to work together. This is similar to engineering ML-enabled systems, whose components have tight coupling (Wan et al., 2019). Participants tried to address this challenge by allowing APIs, such as LangChain (P4, P9, P10) or LangGraph (P19) to handle orchestration, as noted in prior work (Wu et al., 2022a; Dolata et al., 2024).
5.1.3. Implementation
@@
Prompt programming involves constant “experimentation [compared to] traditional software development” (P15) (see Figure 3) and was described as a “trial and error” (P3, P7, P10, P15) process by participants.
Prompt programming is rapid and unsystematic.
@@
Many participants (P1, P4, P6, P7, P8, P12, P13, P14, P15, P19, P20) described a process of starting with a basic prompt for the task to test its feasibility. This prompt included a description of the task and a description of how a human would intuitively perform the task. Each update could result in unwanted behavior and necessitate prompt updates, thus going around the update-test loop again. The prompt grew in complexity as the programmer updated the prompt and observed the model’s output; this was done in tandem with building a mental model of the FM (see Section 4.1.1). This process was “unscientific” (P4): “I would try a lot of different things on a prompt and it’s not a very well thought-out procedure” (P6). It also occurred when participants extended existing prompt programs (P14). Prior work has characterized prompt engineering with FMs as highly iterative (Parnin et al., 2023; Zamfirescu-Pereira et al., 2023a, b; Arawjo et al., 2023; Zamfirescu-Pereira et al., 2023c; Dolata et al., 2024; Lin, 2024; Liu et al., 2024), similar to the experimental nature of engineering ML-enabled systems (Giray, 2021; Nahar et al., 2022, 2023; Wan et al., 2019; Wolf and Paine, 2020).
@@
During this experimentation process, “the speed and the velocity of making changes [was] high” (P19). Literature indicates that exploring the input space for an optimal prompt is challenging (Mishra et al., 2023; Arawjo et al., 2023) and the rapid experimentation with FMs produces many outputs, which are difficult to make sense of (Jiang et al., 2022, 2023; Gero et al., 2022; Liu et al., 2024). Participants created many versions of the prompt to test new prompting approaches (P1, P5, P19), add new functionalities (P2, P6, P13, P16), try to fix bugs (all participants), experiment with new models or hyperparameters (P4), or incorporate feedback after a code review (P1, P14, P16). Prompt versions often branched off from each other (P4, P6, P12, P13, P15, P17, P19, P20), but could be reverted (P4, P5, P8, P10, P15, P16, P20).
@@
As noted by earlier studies (Zamfirescu-Pereira et al., 2023b), changes were less drastic as the prompt stabilized and as the modifications had less of an effect on the prompt, leading “prompt engineering [to feel] like a waste of time” (P5): “If we graphed the number of…changes over the number of iterations, the gradient would be pretty high during the first iterations. Then it flattens quite a bit…because the prompt has matured enough…such that we don’t need to make those big changes anymore” (P15). Thus, one participant avoided testing small prompt changes, and instead preferred to batch them together to reduce costs: “For small changes, we will wait and merge this evaluation into the next time we conduct it” (P16).
@@
Consistent with prior work in prompt engineering (Parnin et al., 2023; Jiang et al., 2022) and building ML-enabled systems (Nahar et al., 2022, 2023; Arpteg et al., 2018; Giray, 2021), the AI’s nondeterminism makes it “hard to measure overall progress” (P6) on the prompt: “You can think you have something that’s really good, and then try it out on a completely different set of inputs and it’s just terrible. A lot of this is feeling of making progress and then backtracking” (P10). Participants developed strategies to address this challenge, like making small changes (P2, P4, P8, P15) or viewing evaluation metrics computed across a large number of examples (P4, P18): “I feel like I wasn’t that confident, other than the fact that our intrinsic numbers went up” (P4).
@@
Participants used a variety of tools to implement their prompt. GUI-based tools (P2, P4, P5, P6, P8) (e.g., GPT playground) and notebooks (P1, P3, P5, P18, P19) enabled rapid prototyping; text editors and IDEs (P1, P11, P14, P17, P20) revealed formatting, like newline characters; and FMs helped inspire or generate parts of the prompt (P2, P18). Prompts were then transferred between environments, usually through copy-pasting (P1, P2, P3, P4, P5, P8). This manual transfer at times introduced issues, such as formatting errors (P1, P5) and inconsistent performance between the model on the external environment compared to the code environment (P5).
@@
Due to the rapid experimentation, participants (P3, P18, P19) struggled to recall previous versions and their outputs. Some tried to address this issue by tracking versions of the prompt as found in Parnin et al. (2023)’s study. This included existing tools such as Git (P5, P6, P7, P10, P16, P17, P18). Other participants set up their own systems to save prompt versions and their performance. This varied in sophistication, from a text document tracking all prompts and outputs (P1, P2, P5, P12) and arbitrary file-naming systems (P9, P12, P19), to spreadsheet tables (P1, P4, P5) and custom-built tools (P8). However, participants still found it difficult to retrieve previous prompt versions (P4, P12) and confused versions due to subtle differences between prompt versions (P9, P12).
5.1.4. Debugging
Prompts can have defects or undesired outputs, which requires debugging from the programmer. Our participants observed a range of defects, such as hallucinations (P9, P14, P15, P17, P19) or not following directions (P2, P3, P4, P6, P8, P9, P10, P12, P18, P19, P20).
Fault localization is not certain.
@@ 5.1.4. Debugging
To fix errors, participants engaged in debugging behaviors. However, debugging prompts is challenging (Jiang et al., 2022; Dolata et al., 2024). The stochastic and black-box nature of FMs made it impossible to ascertain the defect’s source: “Coming from a software engineering background, you…want to set breakpoints and debug, looking at the results step by step. There’s no such mechanism for prompts” (P13). Thus, one challenge was knowing how to change the prompt to address the defect (Mishra et al., 2023): “I don’t even know what to change in my prompt to get there” (P5). Thus, participants engaged in shotgun debugging by trying random changes to the prompt to fix the defect (P2, P7, P8, P10, P15, P18): “In most cases, it is like hit and trial. So maybe I will try a different prompt. I will try it with a different example” (P18). Prior experience and intuition were thus helpful in debugging (P2, P4, P15, P16). Participants also described strategies to gain more confidence in understanding what could fix the error. This included having the FM generate reasoning (P1, P3, P9, P13, P19); testing small, incremental changes (P12); and restricting instructions to specific parts of the prompt (P14, P20). However, participants reported that fixes did not always fix the error (P3, P4, P5, P6, P15, P16, P18, P20), as it could introduce new errors (P3, P4, P7) as documented in prior work (Parnin et al., 2023).
5.1.5. Data curation
@@
To develop prompt programs, programmers must develop new, high-quality datasets, a process referred to as “data curation” (Qian et al., 2024). This became a central activity.
Programmers need to find representative data for the task.
@@
The shift to FM-based components put greater emphasis on finding high-quality data, consistent with prior work (Hassan et al., 2024; Nahar et al., 2022; Arpteg et al., 2018; Dolata and Crowston, 2023; Giray, 2021; Wan et al., 2019; Nahar et al., 2023). To evaluate the prompt (see Section 5.1.6) and provide examples for few-shot learning (Brown, 2020), data curation became important. Since many of the participants’ tasks were custom or specific, using existing benchmarks was not sufficient except in general and common tasks such as jailbreaking (P19). Participants created their own datasets by mining data from the internet or from the participant’s organization, as well as annotating data (P1, P3, P4, P5, P6, P7, P8, P11, P13, P14, P15, P17, P18, P20). The challenge was finding representative data for the task that worked in practice, as prompt programmers needed “to do [their] best to predict how people are going to use [the application]” (P20). Datasets did not necessarily work in practice due to training-serving skew (i.e., when training data does not generalize to production data) (Nahar et al., 2022, 2023; Wan et al., 2019): “You can test with your own dataset, but at the end of the day, you’re still not completely sure how good the output is” (P7).
@@
Participants took care in capturing the full diversity of the input distribution by dividing the inputs into categories and finding examples for each category (P1, P8, P9, P15, P19, P20). Prior work indicates the non-determinism of ML models could pose difficulties in evaluation (Wan et al., 2019; Hassan et al., 2024)—participants addressed this by paraphrasing the inputs (P16, P18, P19, P20). Despite their best efforts, the datasets participants created did not capture the full distribution of inputs. Participants relied on several methods to obtain more data, such as by using an FM to generate examples (P12, P18, P19); having colleagues stress test, “data bash” (P13), or red team (Furmakiewicz et al., 2024) the prompt (P7, P10, P13, P17, P19); and deploying the prompt for user feedback (P7, P10, P13, P15, P17, P19).
@@
Participants struggled with creating representative datasets, as found in prior studies (Arpteg et al., 2018). Despite the emphasis on safety and fairness in literature (Giray, 2021; Arpteg et al., 2018), the prompt program work less well on culturally diverse audiences upon deployment (P7, P10, P13). This reflects prior research in NLP, which found that FMs and datasets align with certain demographics (i.e., Western, college-educated, and younger populations) more than others (Santy et al., 2023; Pei and Jurgens, 2023; Joshi et al., 2020). Additionally, participants found aspects of manual data annotation challenging. Some participants (P3, P13) described challenges on determining labels for the dataset. Participants noted instances of low-quality data annotations: “We found that the [annotations from non-developer colleagues] would be incorrect… It is possible that the worldview of the person doing the annotation is not scoped…so they have some limitations on how much they know [so] they don’t know what the expected outcome is” (P13). Additionally, it the correct label for data was sometimes was unclear: “When do you classify as [Label A]? When do you classify as [Label B]? It was fuzzy on how you would interpret [the labels]” (P3). This reflects documented challenges of handling annotator disagreement and subjectivity (Aroyo and Welty, 2015) in NLP research, where for many tasks, the ground truth can be unclear.
5.1.6. Evaluation
Prompt programmers engage in frequent evaluation of the prompt program.
Evaluating prompt programs requires assessing qualitative constructs.

### v2

@@ 5. Prompt Programming Process
The prompt programming process has significant differences compared to standard software development (see Table 4.3). We identified six activities in prompt programming from axial coding: requirements (Section 5.1), design (Section 5.2), implementation (Section 5.3), debugging (Section 5.4, data curation (Section 5.5), and evaluation (Section 5.6).
5.1. Requirements
@@
In traditional software, requirements are separate from code, while in prompt programming, they are stated in the prompt and may evolve with the behavior of the FM. As in prior work (Furmakiewicz et al., 2024; Parnin et al., 2023), participants noted functional and non-functional requirements, such as usability (P13, P15), reliability (P7, P19), latency (P4, P11, P15, P19), safety (P7, P10, P13, P19), privacy (P4, P13, P14), availability (P4), and cost (P5, P8, P10, P11, P14, P16, P19, P20). Companies found safety and privacy important, as noted in the literature (Hassan et al., 2024; Patel et al., 2025).
[…]
@@
At times, the FM performed an incorrect action not because of an FM error, but due to an under-specified or conflicting requirement in the prompt (P3, P4, P5, P6, P8, P12, P19). Thus, prompt programming required explicitly stating assumptions: “When you’re interacting with a person, you make all these assumptions, but everybody can read between the lines… The more you design prompts, the more you realize that it’s not the same as communicating with another person. The model needs more specific instructions” (P3). However, predicting which assumptions to elaborate on was challenging, as “it’s hard to tell when the model could go wrong” (P3): “I think observation is like 80% [of the requirements]. I knew what I wanted 20% of the time” (P8). Thus, “capturing all the rules…is very difficult in the start, and just generalizing over all the use cases can be very difficult” (P19).
[…]
@@
Similar to building ML-enabled systems (Wan et al., 2019), prompt requirements could change as the FM’s capability on the task was discovered. This included making new trade-offs between non-functional requirements, such as safety vs. quality (Furmakiewicz et al., 2024) (P7): “Our requirements did change as to how the product was perceived… So, striving for that balance between the response quality and the latency induced quite a few iterations” (P15). The participants also defined new features after observing how the FM performed on the task (P2, P4, P13, P18, P20): “I do add in requirements as I iterate. ‘Oh, I didn’t know that you would be able to do this. Let’s see how much I could push this functionality”’ (P2).
5.2. Design
Just as traditional software could be decomposed to base components of classes and functions, complex prompts could be decomposed into multiple tightly coupled prompts.
[…]
@@ 5.2. Design
Participants decomposed prompt programs into smaller components. For a single prompt, the participants decomposed the prompt into subcomponents such as subsections (P1, P2, P5, P6, P7, P8, P13, P20). One participant noted “it is important to structure the prompt in a way that makes sense to you. Otherwise, you’re not able to maintain it going forward” (P13).
@@
Prompts can be composed or “chained” together for greater capabilities (Liu et al., 2023; Wu et al., 2022a, b; Arawjo et al., 2023). The participants decided how to decompose the prompts: “We know what the complex task looks like, and then our job is to break that down” (P9). Each prompt had its own responsibility: “If you are creating a robot, the motion of the hand in any scenario could be a responsibility [and] the movement of the legs could be another” (P19). Similarly to pure code systems, prompts acted as a single module with its own concern (Dijkstra, 1982) that could be composed into a more complex system.
[…]
As observed in other studies (Wu et al., 2022a; Dolata et al., 2024), dependent prompts posed challenges since developers needed to form mental models of how the prompts performed together, rather than individually: “Two changes that do really well separately could merge and become a terrible experience for unexplainable reasons” (P13). This is similar to engineering ML-enabled systems, whose components have tight coupling (Wan et al., 2019). Participants (P8, P12, P13, P19) found that integration resulted in breakage and required significant prompt changes to work with each other. When working individually, P19 performed regular integration testing of multiple prompts: “[Integration testing is] highly correlated with how it’s going to perform in a production environment. In the end, we are going to integrate everything” (P19). However, collaboration could pose challenges, since programmers may be unable to test with other in-progress prompts. P13 reported changing how she would “split up the work” (P13) with her colleagues and felt a “rush to get my PR [(pull request)] in before the other one does” (P13) to avoid the additional work associated with integration. Participants tried to address this challenge by allowing APIs like LangChain (P4, P9, P10) or LangGraph (P19) to handle orchestration, as noted in previous work (Wu et al., 2022a; Dolata et al., 2024).
5.3. Implementation
@@
In prompt programming, natural language prompts serve as both the specification and implementation, while traditional software implements specifications in code. As mentioned in Section 4, prompt programming involves constant “experimentation [compared to] traditional software development” (P15) and is a process of “trial and error” (P3, P7, P10, P15).
[…]
@@
To test the prompt’s feasibility, participants (P1, P4, P6, P7, P8, P12, P13, P14, P15, P19, P20) started with a basic prompt for the task which included a description of the task and how a human would intuitively perform the task. Each update could result in unwanted behavior and require prompt updates, thus restarting the update-test loop. While developing a mental model of the FM (Section 4.1), the prompt became more complex as the programmer refined it and observed the FM’s behavior. This process was “unscientific” (P4): “I would try a lot of different things on a prompt and it’s not a very well thought-out procedure” (P6). It also occurred even with existing prompt programs (P14). Previous work has characterized prompt engineering as highly iterative (Parnin et al., 2023; Zamfirescu-Pereira et al., 2023a, b; Arawjo et al., 2023; Zamfirescu-Pereira et al., 2023c; Dolata et al., 2024; Lin, 2024; Liu et al., 2024), similar to the experimental nature of engineering ML-enabled systems (Giray, 2021; Nahar et al., 2022, 2023; Wan et al., 2019; Wolf and Paine, 2020).
@@
During experimentation, “the speed and the velocity of making changes [was] high” (P19). The literature indicates that finding an optimal prompt is challenging (Mishra et al., 2025; Arawjo et al., 2023), as the rapid experimentation produces many outputs that are difficult to make sense of (Jiang et al., 2022, 2023; Gero et al., 2022; Liu et al., 2024). Participants created many prompt versions to test new approaches (P1, P5, P19), add functionality (P2, P6, P13, P16), fix bugs (all participants), experiment with new models or hyperparameters (P4), or incorporate feedback after a code review (P1, P14, P16). Prompt versions often branched off from each other (P4, P6, P12, P13, P15, P17, P19, P20), but could be reverted (P4, P5, P8, P10, P15, P16, P20).
@@
As noted in earlier studies (Zamfirescu-Pereira et al., 2023b), changes were less drastic as the modifications had less of an effect on the prompt, leading “prompt engineering [to feel] like a waste of time” (P5): “If we graphed the number of…changes over the number of iterations, the gradient would be pretty high during the first iterations. Then it flattens quite a bit…because the prompt has matured…such that we don’t need…big changes” (P15). P16 avoided testing small changes and preferred to batch them to reduce costs: “For small changes, we will wait and merge this evaluation into the next time we conduct it” (P16).
@@
Consistent with prior studies in prompt engineering (Parnin et al., 2023; Jiang et al., 2022) and building ML-enabled systems (Nahar et al., 2022, 2023; Arpteg et al., 2018; Giray, 2021), the AI’s nondeterminism makes it “hard to measure overall progress” (P6): “You can think you have something that’s really good, and then try it out on a completely different set of inputs and it’s just terrible. A lot of this is feeling of making progress and then backtracking” (P10). Participants developed strategies to address this, such as making small changes (P2, P4, P8, P15) or viewing the evaluation metrics computed on a large number of examples (P4, P18): “I feel like I wasn’t that confident, other than the fact that our intrinsic numbers went up” (P4).
@@
The study participants used a variety of tools. GUI-based tools (P2, P4, P5, P6, P8) such as GPT playground and notebooks (P1, P3, P5, P18, P19) enabled rapid prototyping; text editors and IDEs (P1, P11, P14, P17, P20) revealed formatting, like newline characters; and FMs helped inspire or generate parts of the prompt (P2, P18). The prompts were manually transferred between environments, usually by copy-pasting (P1, P2, P3, P4, P5, P8). This introduced issues like formatting errors (P1, P5) and inconsistent behavior between using the GUI versus the API environments (P5).
@@
Due to rapid experimentation, the participants (P3, P18, P19) struggled to recall previous versions and their output. Some tried to address this by tracking versions of the prompt as found in Parnin et al. (2023)’s study by using Git (P5, P6, P7, P10, P16, P17, P18) or even custom systems. These systems varied in sophistication, from a text document tracking all prompts and outputs (P1, P2, P5, P12) and arbitrary file-naming systems (P9, P12, P19), to spreadsheet tables (P1, P4, P5) and custom-built tools (P8). However, participants found it difficult to retrieve previous prompt versions (P4, P12) and confused versions due to subtle differences between versions (P9, P12).
5.4. Debugging
Participants observed defects like hallucinations (P9, P14, P15, P17, P19) or not following directions (P2, P3, P4, P6, P8, P9, P10, P12, P18, P19, P20). Unlike traditional software, where debugging tools can isolate faults to a line of code, prompt debugging is unsystematic due to FM stochasticity.
[…]
@@ 5.4. Debugging
To fix errors, the participants debugged prompts, which is known to be challenging (Jiang et al., 2022; Dolata et al., 2024). The stochastic and opaque nature of FMs made it impossible to determine the source of the defect: “Coming from a software engineering background, you…want to set breakpoints…looking at the results step by step. There’s no such mechanism for prompts” (P13). One challenge was to know how to change the prompt to fix the defect (Mishra et al., 2025): “I don’t even know what to change in my prompt to get there” (P5). Participants engaged in shotgun debugging by trying random changes to fix the defect (P2, P7, P8, P10, P15, P18): “In most cases, it is like hit and trial. So maybe I will try a different prompt…[or] example” (P18). Thus, prior experience and intuition were helpful in debugging (P2, P4, P15, P16). Study participants also described strategies to gain more confidence in understanding what could fix the error. This included having the FM generate reasoning (P1, P3, P9, P13, P19); testing small changes (P12); and restricting instructions to specific parts of the prompt (P14, P20). However, the change did not always fix the error (P3, P4, P5, P6, P15, P16, P18, P20), as it could introduce new errors (P3, P4, P7), as documented in previous work (Parnin et al., 2023).
5.5. Data Curation
@@
To develop prompt programs, programmers must create datasets, a process referred to as “data curation” (Qian et al., 2024). This became a central activity that is not present in traditional software engineering.
[…]
@@
Using FM-based components shifts the focus to finding high-quality data, consistent with previous work (Hassan et al., 2024; Nahar et al., 2022; Arpteg et al., 2018; Dolata and Crowston, 2023; Giray, 2021; Wan et al., 2019; Nahar et al., 2023). To evaluate the prompt and provide examples for few-shot learning (Brown et al., 2020), data curation was paramount. Since many participant tasks were custom or specific, the existing benchmarks were not sufficient except for more common tasks like jailbreaking (P19). Study participants created their own datasets by mining data from the internet or from their organization, as well as annotating data (P1, P3, P4, P5, P6, P7, P8, P11, P13, P14, P15, P17, P18, P20). The challenge was to find representative data that worked in practice, as prompt programmers needed “to predict how people are going to use [the application]” (P20). Datasets did not always work in practice due to training-serving skew (i.e., when training data does not generalize to production data) (Nahar et al., 2022, 2023; Wan et al., 2019): “You can test with your own dataset, but at the end of the day, you’re still not completely sure how good the output is” (P7).
@@
The participants ensured diverse inputs by categorizing the inputs and selecting examples for each category (P1, P8, P9, P15, P19, P20). Previous work indicates that the nondeterminism of ML models poses difficulties in evaluation (Wan et al., 2019; Hassan et al., 2024). The participants addressed this by paraphrasing the inputs (P16, P18, P19, P20). Despite their efforts, the datasets did not capture the input distribution, so the participants relied on several methods to obtain more data like using FMs to generate examples (P12, P18, P19); asking colleagues to stress test, “data bash” (P13), or red team (Furmakiewicz et al., 2024) the prompt (P7, P10, P13, P17, P19); and obtaining user feedback (P7, P10, P13, P15, P17, P19).
@@
As found in other work (Arpteg et al., 2018), participants struggled to create datasets. Despite the emphasis on safety and fairness in the literature (Giray, 2021; Arpteg et al., 2018), prompt programs performed worse on culturally diverse audiences during deployment (P7, P10, P13). This reflects NLP research, where FMs and datasets align with certain demographics (i.e., Western, college-educated, and younger populations) more than others (Santy* et al., 2023; Pei and Jurgens, 2023; Joshi et al., 2020). Data labeling was also challenging, as annotations could be of lower quality: “The [annotations from some colleagues] would be incorrect…they have some limitations on how much they know, [so] they don’t know what the expected outcome is” (P13). Meanwhile, the correct label could be unclear (P3, P13): “When do you classify as [Label A]? It was fuzzy on how you would interpret [the labels]” (P3). This reflects the challenges of handling annotator disagreement and subjectivity in NLP research (Aroyo and Welty, 2015), where ground truth labels can be subjective.
5.6. Evaluation
Both traditional software and prompt programs involve testing at multiple scopes. Traditional software uses code coverage, while prompt programming often relies on qualitative evaluation.
[…]